OpenStudy (anonymous):

PLEASE HELP!! solve the equation (e^x-81e^-x)/2 =9 which is the solution in terms of natural logarithms? x=__________ OR there is no solution what is the decimal approximation for the solution? x=______________ OR There is no solution

OpenStudy (anonymous):

\[ e^x-81e^{-x}=18\\ e^x-{81\over e^x}=18\\ e^{2x}-81=18e^x\\ e^{2x}-18e^x-81=0 \] substitute \(u=e^x\) solve for "u" then find "x"

OpenStudy (anonymous):

would it be 9?

OpenStudy (anonymous):

would what be 9?

OpenStudy (anonymous):

u=9

OpenStudy (anonymous):

how will you do the e^x=9

OpenStudy (anonymous):

if the question is correct, there'd be no solutions! check by pluggin in "u=9"

OpenStudy (anonymous):

check if the second e^-x is positive or negative...

OpenStudy (anonymous):

9 dosent work

jhonyy9 (jhonyy9):

so u=9 not is right because there is -81 and not +81 inside quadratic equation

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

check the question

OpenStudy (anonymous):

yes I just saw that.

OpenStudy (anonymous):

saw what? did you have the question wrong?

OpenStudy (anonymous):

lol no saw that 9 doesnt work

OpenStudy (anonymous):

yep. so, no REAL solutions exist.

OpenStudy (anonymous):

okay thank you very much!!

OpenStudy (anonymous):

yer welcome very much.